OpenStudy (anonymous):

HELP FAST I WILL FAN AND MEDAL A number cube rolled. what is P(3 or even number)?

OpenStudy (anonymous):

al see what i can do

OpenStudy (xguardians):

1/6 of being P(3) + 3/6 of being an even number = 4/6

OpenStudy (anonymous):

HOLD ON I STILL HAVE TO PUT THE ANSWERS

OpenStudy (xguardians):

So simplified the answer is 2/3

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

\[A. \frac{ 1 }{ 2 }\] \[B. \frac{ 2 }{ 3 }\] \[C. \frac{ 4 }{ 2 }\] \[D. \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

prob 2/3

OpenStudy (xguardians):

So simplified the answer is 2/3 B. Please read my explanation too.

OpenStudy (anonymous):

i read it

OpenStudy (anonymous):

and @lilcece we cool

OpenStudy (anonymous):

are we cool?

OpenStudy (anonymous):

uhhhhh sure :)

OpenStudy (anonymous):

:3